OpenStudy (anonymous):

can some one explain this question i really appreciated. i am stuck Solve. x2 + 6x + 4 = 0

OpenStudy (lexipoo1998):

It just wants you to solve for x. so isolate the x's and solve.

OpenStudy (therealmeeeee):

ok x² - 6x - 4 = 0 x² - 6x = 4 x² - 6x + 9 = 4 + 9 ( x - 3 ) ( x - 3 ) = 13 x - 3 = ±sqrt ( 13 ) x = 3 ± sqrt ( 13 )

OpenStudy (therealmeeeee):

ok So the answer is 3 +- sqrt 13.

OpenStudy (lexipoo1998):

ya like what he did ^

OpenStudy (therealmeeeee):

do this help

OpenStudy (therealmeeeee):

click Best Response

OpenStudy (anonymous):

@yulima , or u can using this formula: \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

like how i just replace the number in the letters

OpenStudy (therealmeeeee):

Yes:)

OpenStudy (anonymous):

what about the cero

OpenStudy (therealmeeeee):

well u see the answer u can figure out wat happend to it

OpenStudy (anonymous):

 this is what i got